# A $2$-distance $(2\Delta+7)$-coloring of planar graphs

### v2

@@ 2.1. The case $\Delta=6$
Let $f$ be a $5$-face. Suppose first that $f$ is not incident to any $3$-vertex. Note that if $f$ is not incident to any bad $4$-vertices, then $\mu^{*}(v)\geq 1-5\times\frac{1}{5}=0$ after $f$ sends at most $\frac{1}{5}$ to each of its incident vertices by R2(b),(c). Therefore we further suppose that $f$ is incident to at least one bad $4$-vertex, let $u$ be such a bad $4$-vertex incident to $f$. Recall that $u$ has no 6(6)666(6)6 ( 6 )-neighbour by Lemma 2.3. If $u$ is a $4(1)$-vertex, then it has at most one neighbour consisting of bad $4$-vertex or $5(4)$-vertex by Lemma 2.4 and Corollary 2.9. Besides, if $u$ is a $4(2)$-vertex, then the neighbours of $u$ are different from $4$- and $5(4)$-vertices by Lemma 2.7. Consequently, if $f$ is incident to three bad $4$-vertices, then the other vertices incident to $f$ are different from $4$-, $5(4)$- and 6(6)666(6)6 ( 6 )-vertices. Moreover, if $f$ is incident to exactly two bad $4$-vertices, then $f$ is incident to at most one vertex consisting of $5(4)$- or 6(6)666(6)6 ( 6 )-vertices. On the other hand, if $f$ is incident to exactly one bad $4$-vertex $u$, then $u$ has no 6(6)666(6)6 ( 6 )-neighbour, and has at most one $5(4)$-neighbour, so $f$ is incident to at most three vertices consisting of $5(4)$- or 6(6)666(6)6 ( 6 )-vertices. We then conclude that $\mu^{*}(f)\geq 0$, since the face $f$ sends $\frac{1}{3}$ to each of its incident bad $4$-vertices by R2(a), $\frac{1}{5}$ to each of its incident $5(4)$-vertices by R2(b) and $\frac{1}{5}$ to each of its incident 6(6)666(6)6 ( 6 )-vertices having no $4$-neighbour by R2(c).
Now we suppose that $f$ is incident to a $3$-vertex. By Lemma 2.2, $f$ can be incident to at most two $3$-vertices. If $f$ is incident to exactly one $3$-vertex, say $v_{1}$, then the neighbours of $v_{1}$ lying on $f$, denote by $v_{2},v_{5}$, are different from $5^{-}$- and $6(5^{+})$-vertex by Lemma 2.2. Thus, $\mu^{*}(v)\geq 1-3\times\frac{1}{3}=0$ after $f$ sends at most $\frac{1}{3}$ to each of its incident vertices other than $v_{2},v_{5}$ by R2. Next we assume that $f$ is incident to exactly two $3$-vertices, say $v_{1},v_{3}$, then all vertices incident to $f$ other than $v_{1},v_{3}$ are different from $5^{-}$- and $6(5^{+})$-vertex by Lemma 2.2. Thus, $\mu^{*}(v)\geq 1-2\times\frac{1}{3}>0$ after $f$ sends $\frac{1}{3}$ to each of $v_{1},v_{3}$ by R2(a).
@@
(1). Let $k=3$. By Lemma 2.2, $m_{3}(v)=0$, $m_{4}(v)\leq 1$, and each neighbour of $v$ is a $6(4^{-})$-vertex. Note that $v$ is incident to at least two $5^{+}$-faces, and each of them gives $\frac{1}{3}$ to $v$ by R2(a). On the other hand, $v$ is incident to three $6(4^{-})$-vertices, and each of them gives at least $\frac{1}{9}$ to $v$ by R3 and R4(b). Thus, we have $\mu^{*}(v)\geq-1+2\times\frac{1}{3}+3\times\frac{1}{9}\geq 0$.
(2). Let $k=4$. By Lemma 2.3, $v$ is incident to at most two $3$-faces, in particular, $v$ has no 6(6)666(6)6 ( 6 )-neighbour. Clearly, we have $\mu^{*}(v)\geq 0$ when $v$ is not incident to any $3$-face, since $v$ does not give a charge to any vertex or face. We may therefore assume that $1\leq m_{3}(v)\leq 2$; namely, $v$ is a bad $4$-vertex.
First suppose that $m_{3}(v)=1$. Let $m_{4}(v)=r$ for $0\leq r\leq 3$. By Lemma 2.5, $n_{6}(v)\geq r+1$, i.e., $v$ has at least $r+1$ many $6(5^{-})$-neighbours, and each of them gives at least $\frac{1}{12}$ to $v$ by R3, R4(a) and R5(a). On the other hand, $v$ is incident to $3-r$ many $5^{+}$-faces, and each of them gives $\frac{1}{3}$ to $v$ by R2(a). Consequently, $v$ receives totally $(3-r)\times\frac{1}{3}$ from its incident $5^{+}$-faces, and totally $(r+1)\times\frac{1}{12}$ from its $6$-neighbours. Thus, $\mu^{*}(v)\geq(3-r)\times\frac{1}{3}+(r+1)\times\frac{1}{12}-\frac{1}{3}\geq 0$ after $v$ sends $\frac{1}{3}$ to its incident $3$-face by R1.
We now suppose that $m_{3}(v)=2$, and let $m_{4}(v)=r$. Then, by Lemma 2.6, we have $0\leq r\leq 1$ and $n_{6}(v)\geq r+3$. Since $v$ is not adjacent to any 6(6)666(6)6 ( 6 )-vertex, $v$ has at least $r+3$ many $6(5^{-})$-neighbours, and each of them gives at least $\frac{1}{12}$ to $v$ by R3, R4(a) and R5(a). On the other hand, $v$ is incident to $2-r$ many $5^{+}$-faces, and each of them gives $\frac{1}{3}$ to $v$ by R2(a). Consequently, $v$ receives totally $(2-r)\times\frac{1}{3}$ from its incident $5^{+}$-faces, and totally $(r+3)\times\frac{1}{12}$ from its $6$-neighbours. Thus, $\mu^{*}(v)\geq(2-r)\times\frac{1}{3}+(r+3)\times\frac{1}{12}-2\times\frac{1}{3}\geq 0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1.
@@
Let $m_{3}(v)=4$. If $m_{4}(v)=0$, then $v$ has at least two $6(5^{-})$-neighbours by Lemma 2.10(a), and $v$ receives $\frac{1}{9}$ from each of its $6(5^{-})$-neighbours by R3, R4(b) and R5(b). Also, $v$ receives $\frac{1}{5}$ from its incident $5^{+}$-face by R2(b). Thus, $\mu^{*}(v)\geq 1+2\times\frac{1}{9}+\frac{1}{5}-4\times\frac{1}{3}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1. Suppose now that $m_{4}(v)=1$. Then $v$ has at least three $6(5^{-})$-neighbours by Lemma 2.10(b). It follows from applying R3, R4(b) and R5(b) that $v$ receives at least $\frac{1}{9}$ from each of its $6(5^{-})$-neighbours. Thus, $\mu^{*}(v)\geq 1+3\times\frac{1}{9}-4\times\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1.
Let $m_{3}(v)=5$. Recall that $v$ has no 6(6)666(6)6 ( 6 )-neighbour by Lemma 2.12. Suppose first that there exists $uw\in G[N(v)]$ such that $uw$ is contained in two $3$-faces. It then follows from Proposition 2.14(a) that all neighbours of $v$ are $6(5^{-})$-vertices, and each $vv_{i}$ for $v_{i}\in N(v)\setminus\{u,w\}$ is a special edge. This particularly implies that each $v_{i}\in N(v)\setminus\{u,w\}$ is a $6(4^{-})$-vertex. Then, $v$ receives $\frac{1}{6}$ from each $v_{i}$ for $i\in[3]$ by R3 and R4(c), and $\frac{1}{9}$ from each of $u,w$ by R4(d) and R5(b). Thus, $v$ totally receives at least $\frac{2}{3}$ from its $6$-neighbours. Next we suppose that no edge $uw\in G[N(v)]$ is contained in two $3$-faces. By Proposition 2.14(b), there exist four $6(4^{-})$-vertices $v_{1},v_{2},v_{3},v_{4}\in N(v)$ such that each $vv_{i}$ is a special edge. Then, $v$ receives $\frac{1}{6}$ from each $v_{i}$ for $i\in[4]$ by R3 and R4(c). Clearly, $v$ receives totally at least $\frac{2}{3}$ from its $6$-neighbours. Hence, $\mu^{*}(v)\geq 1+\frac{2}{3}-5\times\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1.
@@
(4.1). Let $m_{3}(v)=4$. Obviously, $v$ has one of the three configurations depicted in Figure 2. Notice that $v$ only gives charge to its incident $3$-faces, and $4^{-}$-, $5(4^{+})$-, $6(5)$-neighbours by R1, R4. We first note that if $v$ has no 5(5)555(5)5 ( 5 )-neighbour $v_{i}$ such that $vv_{i}$ is a special edge, then $v$ gives at most $\frac{1}{9}$ to each of its neighbours by R4(a),(b),(d), and so $\mu^{*}(v)\geq 2-6\times\frac{1}{9}-4\times\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1. We may further assume that $v$ has at least one 5(5)555(5)5 ( 5 )-neighbour $v_{i}$ such that $vv_{i}$ is a special edge. In fact, we will determine the final charge of $v$ based on the number of 5(5)555(5)5 ( 5 )-neighbors $v_{i}$ of $v$ for which each edge $vv_{i}$ is a special edge. It can be easily observed that $v$ cannot have three 5(5)555(5)5 ( 5 )-neighbours by Lemma 2.12.
$v$$v_{3}$$v_{4}$$v_{5}$$v_{6}$$v_{1}$$v_{2}$
$v$$v_{3}$$v_{4}$$v_{5}$$v_{6}$$v_{1}$$v_{2}$
$v$$v_{3}$$v_{4}$$v_{5}$$v_{6}$$v_{1}$$v_{2}$
@@
Suppose first that $v$ has two 5(5)555(5)5 ( 5 )-neighbours $v_{i}$ and $v_{j}$ such that both $vv_{i}$ and $vv_{j}$ are special edges. By Lemma 2.12, $v_{i}$ and $v_{j}$ are non-adjacent, so the configuration in Figure 2(b) is not possible for $v$. Moreover, $v$ cannot have the form in Figure 2(c) as well, since $vv_{i}$ and $vv_{j}$ are special edges. Hence, we deduce that the neighbours of $v$ can only form as depicted in Figure 2(a), where $v_{i}=v_{2}$ and $v_{j}=v_{5}$. By Lemmas 2.2 and 2.12, $v_{2}$ and $v_{5}$ cannot have a $4^{-}$-neighbour, bad $5$-neighbour and 6(6)666(6)6 ( 6 )-neighbour, so the other neighbours of $v$ are $6(5^{-})$-vertices or $5(t)$-vertices for $0\leq t\leq 3$. Thus, $v$ gives $\frac{1}{6}$ to each of $v_{2},v_{5}$ by R4(c), and at most $\frac{1}{12}$ to each of the other neighbours of $v$ by R4(a). Consequently, $\mu^{*}(v)\geq 2-2\times\frac{1}{6}-4\times\frac{1}{12}-4\times\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1.
Now we suppose that $v$ has exactly one 5(5)555(5)5 ( 5 )-neighbour $v_{i}$ such that $vv_{i}$ is a special edge. Then, similarly as above, we deduce that the neighbours of $v$ can only form as depicted in Figure 2(a) where $v_{i}=v_{2}$ (or $v_{i}=v_{5}$). By Lemmas 2.2 and 2.12, $v_{2}$ cannot have a $4^{-}$-neighbour, bad $5$-neighbour and 6(6)666(6)6 ( 6 )-neighbour, so the neighbours of $v_{2}$ other than $v$ are $6(5^{-})$-vertices or $5(t)$-vertices for $0\leq t\leq 3$. Thus, $v$ gives $\frac{1}{6}$ to $v_{2}$, at most $\frac{1}{12}$ to each of $v_{1},v_{3}$ by R4(a), at most $\frac{1}{9}$ to $v_{5}$ by R4(a),(b),(d), and at most $\frac{1}{12}$ to each of $v_{4},v_{6}$ by R4(a),(b),(d). Hence, $\mu^{*}(v)\geq 2-\frac{1}{6}-2\times\frac{1}{12}-\frac{1}{9}-2\times\frac{1}{12}-4\times\frac{1}{3}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1.
@@
Let $n_{4}(v)=1$. By Lemma 2.15(d), $v$ has at most three bad $5$-neighbours. If $v$ has such three neighbours, then $v$ would have a $6(4^{-})$-neighbour by Lemma 2.15(e), and so $v$ receives at least $\frac{1}{12}$ from its $6(4^{-})$-neighbour by R3 and R4(a). On the other hand, $v$ gives $\frac{1}{12}$ to its bad $4$-neighbour by R5(a), and $\frac{1}{9}$ to each of its bad $5$-neighbours by R5(b). Thus, $\mu^{*}(v)\geq 2+\frac{1}{12}-\frac{1}{12}-3\times\frac{1}{9}-5\times\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1. Suppose now that $v$ has at most two bad $5$-vertices. Then, $\mu^{*}(v)\geq 2-\frac{1}{12}-2\times\frac{1}{9}-5\times\frac{1}{3}>0$ after $v$ sends $\frac{1}{12}$ to its bad $4$-neighbour R5(a), $\frac{1}{9}$ to each of its bad $5$-neighbours R5(b), and $\frac{1}{3}$ to each of its incident $3$-faces by R1.

### v3

@@ 2.1. The case $\Delta=6$
Let $f$ be a $5$-face. Suppose first that $f$ is not incident to any $3$-vertex. Note that if $f$ is not incident to any bad $4$-vertices, then $\mu^{*}(v)\geq 1-5\times\frac{1}{5}=0$ after $f$ sends $\frac{1}{5}$ to each of its incident vertices by R2(b),(c). Therefore we further suppose that $f$ is incident to at least one bad $4$-vertex, let $u$ be such a bad $4$-vertex incident to $f$. If $u$ is a $4(1)$-vertex, then it has at most one neighbour consisting of $4$- or $5(4)$-vertex by Lemma 2.6 and Corollary 2.11. Besides, if $u$ is a $4(2)$-vertex, then the neighbours of $u$ are neither $4$-vertex nor $5(4)$-vertex by Lemma 2.9. As a result, $f$ is incident to at most three bad $4$-vertices. On the other hand, a $5(4)$-vertex having a $4$-neighbour cannot have any other $4$- or $5(4)$-neighbour by Lemma 2.10. Thus, we conclude the following.
  • If $f$ is incident to three bad $4$-vertices, then the remaining vertices incident to $f$ are neither a $4$-vertex, nor a $5(4)$-vertex, nor a $6(5)$-vertex having no $4$-neighbour.
  • If $f$ is incident to exactly two bad $4$-vertices, then $f$ is incident to at most one vertex that is either a $5(4)$-vertex or a $6(5)$-vertex having no $4$-neighbour.
  • If $f$ is incident to exactly one bad $4$-vertex, say $w$, then one of the neighbors of $w$ lying on $f$ is neither a $4$-vertex, nor a $5(4)$-vertex, nor a $6(5)$-vertex having no $4$-neighbour.
In each case, we deduce that $\mu^{*}(f)\geq 0$, since the face $f$ sends $\frac{1}{3}$ to each of its incident bad $4$-vertices by R2(a), $\frac{1}{5}$ to each of its incident $5(4)$-vertices by R2(b) and $\frac{1}{5}$ to each of its incident $6(5)$-vertices having no $4$-neighbour by R2(c).
Now we suppose that $f$ is incident to a $3$-vertex, say $v_{1}$. Then the neighbours of $v_{1}$ lying on $f$, denote by $v_{2},v_{5}$, are different from $5^{-}$- and $6(5^{+})$-vertex by Lemma 2.4. Thus, $\mu^{*}(v)\geq 1-3\times\frac{1}{3}=0$ after $f$ sends at most $\frac{1}{3}$ to each of its incident vertices other than $v_{2},v_{5}$ by R2.
@@
(1). Let $k=3$. By Lemma 2.4, $m_{3}(v)=0$, $m_{4}(v)\leq 1$, and each neighbour of $v$ is a $6(4^{-})$-vertex. Note that $v$ is incident to at least two $5^{+}$-faces, and each of them gives $\frac{1}{3}$ to $v$ by R2(a). On the other hand, $v$ is incident to three $6(4^{-})$-vertices, and each of them gives at least $\frac{1}{9}$ to $v$ by R3 and R4(b). Thus, we have $\mu^{*}(v)\geq-1+2\times\frac{1}{3}+3\times\frac{1}{9}=0$.
(2). Let $k=4$. By Lemma 2.5, $v$ is incident to at most two $3$-faces, in particular, $v$ has no $6(6)$-neighbour. Clearly, we have $\mu^{*}(v)\geq 0$ when $v$ is not incident to any $3$-face, since $v$ does not give a charge to any vertex or face. We may therefore assume that $1\leq m_{3}(v)\leq 2$; namely, $v$ is a bad $4$-vertex.
First suppose that $m_{3}(v)=1$. Let $m_{4}(v)=r$ for $0\leq r\leq 3$. By Lemma 2.7, we have $n_{6}(v)\geq r+1$, i.e., $v$ has at least $r+1$ many $6(5^{-})$-neighbours, and each of them gives at least $\frac{1}{12}$ to $v$ by R3, R4(a) and R5(a). On the other hand, $v$ is incident to 3r3-r many $5^{+}$-faces, and each of them gives $\frac{1}{3}$ to $v$ by R2(a). Consequently, $v$ receives totally $(3-r)\times\frac{1}{3}$ from its incident $5^{+}$-faces, and totally $(r+1)\times\frac{1}{12}$ from its $6(5^{-})$-neighbours. Thus, $\mu^{*}(v)\geq(3-r)\times\frac{1}{3}+(r+1)\times\frac{1}{12}-\frac{1}{3}\geq 0$ after $v$ sends $\frac{1}{3}$ to its incident $3$-face by R1.
We now suppose that $m_{3}(v)=2$. Then, by Lemma 2.8, we have $m_{4}(v)=r\leq 1$ and $n_{6}(v)\geq r+3$. Since $v$ is not adjacent to any $6(6)$-vertex, $v$ has at least $r+3$ many $6(5^{-})$-neighbours, and each of them gives at least $\frac{1}{12}$ to $v$ by R3, R4(a) and R5(a). On the other hand, $v$ is incident to 2r2-r many $5^{+}$-faces, and each of them gives $\frac{1}{3}$ to $v$ by R2(a). Consequently, $v$ receives totally $(2-r)\times\frac{1}{3}$ from its incident $5^{+}$-faces, and totally $(r+3)\times\frac{1}{12}$ from its $6$-neighbours. Thus, $\mu^{*}(v)\geq(2-r)\times\frac{1}{3}+(r+3)\times\frac{1}{12}-2\times\frac{1}{3}\geq 0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1.
@@
Let $m_{3}(v)=4$. If $m_{4}(v)=0$, then $v$ has at least two $6(5^{-})$-neighbours by Lemma 2.12(a), and $v$ receives at least $\frac{1}{9}$ from each of its $6(5^{-})$-neighbours by R3, R4(b) and R5(b). Also, $v$ receives $\frac{1}{5}$ from its incident $5^{+}$-face by R2(b). Thus, $\mu^{*}(v)\geq 1+2\times\frac{1}{9}+\frac{1}{5}-4\times\frac{1}{3}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1. Suppose now that $m_{4}(v)=1$. Then $v$ has at least three $6(5^{-})$-neighbours by Lemma 2.12(b). It follows from applying R3, R4(b) and R5(b) that $v$ receives at least $\frac{1}{9}$ from each of its $6(5^{-})$-neighbours. Thus, $\mu^{*}(v)\geq 1+3\times\frac{1}{9}-4\times\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1.
Let $m_{3}(v)=5$. Recall that $v$ has no $6(6)$-neighbour by Lemma 2.14. Suppose first that there exists $v_{i}v_{i+1}\in E(v)$ such that $v_{i}v_{i+1}$ is contained in two $3$-faces. It then follows from Proposition 2.16(a) that all neighbours of $v$ are $6(5^{-})$-vertices, and each $vv_{j}$ for $v_{j}\in N(v)\setminus\{v_{i},v_{i+1}\}$ is a special edge. This particularly implies that each $v_{j}\in N(v)\setminus\{v_{i},v_{i+1}\}$ is a $6(4^{-})$-vertex. Then, $v$ receives $\frac{1}{6}$ from each $v_{j}$ for $j\in[5]\setminus\{i,i+1\}$ by R3 and R4(c), and at least $\frac{1}{9}$ from each of $v_{i},v_{i+1}$ by R3, R4(d) and R5(b). Thus, $v$ totally receives at least $\frac{2}{3}$ from its $6$-neighbours. Next we suppose that no edge $v_{i}v_{i+1}\in E(v)$ is contained in two $3$-faces. By Proposition 2.16(b), there exist four $6(4^{-})$-vertices $v_{1},v_{2},v_{3},v_{4}\in N(v)$ such that each $vv_{i}$ is a special edge. Then, $v$ receives $\frac{1}{6}$ from each $v_{i}$ for $i\in[4]$ by R3 and R4(c). Clearly, $v$ receives totally at least $\frac{2}{3}$ from its $6$-neighbours. Hence, $\mu^{*}(v)\geq 1+\frac{2}{3}-5\times\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1.
@@
(4.1). Let $m_{3}(v)=4$. Obviously, $v$ has one of the three configurations depicted in Figure 2. Notice that $v$ only gives charge to its incident $3$-faces, and $4^{-}$-, $5(4^{+})$-, $6(5)$-neighbours by R1, R4. We first note that if $v$ has no $5(5)$-neighbour $v_{i}$ such that $vv_{i}$ is a special edge, then $v$ gives at most $\frac{1}{9}$ to each of its neighbours by R4(a),(b),(d), and so $\mu^{*}(v)\geq 2-6\times\frac{1}{9}-4\times\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1. We may further assume that $v$ has at least one $5(5)$-neighbour $v_{i}$ such that $vv_{i}$ is a special edge. In fact, we will determine the final charge of $v$ based on the number of $5(5)$-neighbors $v_{i}$ of $v$ for which each edge $vv_{i}$ is a special edge. It can be easily observed that $v$ cannot have three $5(5)$-neighbours by Lemma 2.14.
[…]
@@
Suppose first that $v$ has two $5(5)$-neighbours $v_{i}$ and $v_{j}$ such that both $vv_{i}$ and $vv_{j}$ are special edges. By Lemma 2.14, $v_{i}$ and $v_{j}$ are non-adjacent, so the configuration in Figure 2(b) is not possible for $v$. Moreover, $v$ cannot have the form in Figure 2(c) as well, since $vv_{i}$ and $vv_{j}$ are special edges. Hence, we deduce that the neighbours of $v$ can only form as depicted in Figure 2(a), where $v_{i}=v_{2}$ and $v_{j}=v_{5}$. By Lemmas 2.4 and 2.14, $v_{2}$ and $v_{5}$ cannot have a $4^{-}$-neighbour, bad $5$-neighbour and $6(6)$-neighbour, so the other neighbours of $v$ are $6(5^{-})$-vertices or $5(3^{-})$-vertices. Thus, $v$ gives $\frac{1}{6}$ to each of $v_{2},v_{5}$ by R4(c), and at most $\frac{1}{12}$ to each of the other neighbours of $v$ by R4(a). Consequently, $\mu^{*}(v)\geq 2-2\times\frac{1}{6}-4\times\frac{1}{12}-4\times\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1.
Now we suppose that $v$ has exactly one $5(5)$-neighbour $v_{i}$ such that $vv_{i}$ is a special edge. Then, similarly as above, we deduce that the neighbours of $v$ can only form as depicted in Figure 2(a) where $v_{i}=v_{2}$ (or $v_{i}=v_{5}$). By Lemmas 2.4 and 2.14, $v_{2}$ cannot have a $4^{-}$-neighbour, bad $5$-neighbour and $6(6)$-neighbour, so the neighbours of $v_{2}$ other than $v$ are $6(5^{-})$-vertices or $5(3^{-})$-vertices. Thus, $v$ gives $\frac{1}{6}$ to $v_{2}$, at most $\frac{1}{12}$ to each of $v_{1},v_{3}$ by R4(a), at most $\frac{1}{9}$ to each of $v_{4},v_{5},v_{6}$ by R4(a),(b),(d). Hence, $\mu^{*}(v)\geq 2-\frac{1}{6}-2\times\frac{1}{12}-3\times\frac{1}{9}-4\times\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1.
@@
Let $n_{4}(v)=1$. By Lemma 2.17(d), $v$ has at most three bad $5$-neighbours. If $v$ has such three neighbours, then $v$ has a $6(4^{-})$-neighbour by Lemma 2.17(e), and so $v$ receives at least $\frac{1}{12}$ from its $6(4^{-})$-neighbour by R3 and R4(a). On the other hand, $v$ gives $\frac{1}{12}$ to its bad $4$-neighbour by R5(a), and $\frac{1}{9}$ to each of its bad $5$-neighbours by R5(b). Thus, $\mu^{*}(v)\geq 2+\frac{1}{12}-\frac{1}{12}-3\times\frac{1}{9}-5\times\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1. Suppose now that $v$ has at most two bad $5$-vertices. Then, $\mu^{*}(v)\geq 2-\frac{1}{12}-2\times\frac{1}{9}-5\times\frac{1}{3}>0$ after $v$ sends $\frac{1}{12}$ to its bad $4$-neighbour R5(a), $\frac{1}{9}$ to each of its bad $5$-neighbours R5(b), and $\frac{1}{3}$ to each of its incident $3$-faces by R1.